OpenStudy (anonymous):

Can anyone help me with this? For which \[p \in R \] is the function \[f _{p} (x)= \frac{ 1 }{ (x-1)\log ^{p}_{}(x) } \] improperly Riemann integrable on the interval \[\left[ e,\infty \right] \] ?

OpenStudy (anonymous):

What would you guess be? What definitions apply here? Any relevant theorems?

OpenStudy (anonymous):

for p=0, it is not.

OpenStudy (anonymous):

Could you explain me how?

OpenStudy (anonymous):

Notice that \[ \lim_{x->\infty} \frac {x \ln^p(x) }{(x-1) \ln^p(x)}=1 \]

OpenStudy (anonymous):

Now for \[ \int_e^b \frac{dx}{ x \ln^p(x)}\\ u=\ln(x)\\ du =\frac {dx} x\\ \int_e^b \frac{dx}{ x \ln^p(x)}=\int_1^{\ln(b)}\frac {du}{u^p} \] The last integral converges if \( p>1\) and diverges if \( p\le1\) Hence the original integral do the same by the limit convergence theorem.

OpenStudy (anonymous):

Did you understand the proof?

OpenStudy (anonymous):

yeah..sort of :) thank you so much... :)

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

can you help for this as well ? For which p∈R is the infinite series \[\sum_{k=2}^{\infty} =\frac{ k(-1)^{k} }{ (k-1)^{^{2}}\log ^{^{p}}(k) }\] convergent? conditionally convergent? absolutely convergent?

OpenStudy (anonymous):

The rules specify that you should post it as a new question.

OpenStudy (anonymous):

ohh..right.. sorry about that..